OpenStudy (anonymous):

Find the center and radius of the circle: x^2+2x+y^2+10y=15

OpenStudy (anonymous):

\[x^2 + 2x + y^2 +10y = 15\]

OpenStudy (triciaal):

do you know how to complete the square?

OpenStudy (anonymous):

What does that mean?

OpenStudy (triciaal):

equation of a circle (x-h)^2 + (y-k)^2 = r^2 where the center is (h, k) and the radius, r

OpenStudy (anonymous):

Oh yea! But I don't know what I'm doing in this problem

OpenStudy (triciaal):

think of completing 2 squares

OpenStudy (anonymous):

How would I do that?

OpenStudy (triciaal):

I am restricted on this laptop don't have "draw"

OpenStudy (triciaal):

think of completing 2 squares one with x and one with y combine the constants look at the formula above for the equation of a circle

OpenStudy (anonymous):

Could you type out how solving it would look like?